# R-enum Revisited: Speedup and Extension for Context-Sensitive Repeats and Net Frequencies

### v1

@@ Abstract
A repeat is a substring that occurs at least twice in a string, and is called a maximal repeat if it cannot be extended outwards without reducing its frequency. Nishimoto and Tabei [CPM, 2021] proposed r-enum, an algorithm to enumerate various characteristic substrings, including maximal repeats, in a string $T$ of length $n$ in $O(r)$ words of compressed working space, where $r\leq n$ is the number of runs in the Burrows-Wheeler transform (BWT) of $T$. Given the run-length encoded BWT (RLBWT) of $T$, r-enum runs in $O(n\log\log_{w}(n/r))$ time in addition to the time linear to the number of output strings, where $w=\Theta(\log n)$ is the word size. In this paper, we improve the $O(n\log\log_{w}(n/r))$ term to $O(n)$. We also extend r-enum to compute other context-sensitive repeats such as near-supermaximal repeats (NSMRs) and supermaximal repeats, as well as the context-diversity for every maximal repeat in the same complexities. Furthermore, we study the occurrences that witness NSMRs, which have recently attracted attention under the name of net occurrences: An occurrence of a repeat is called a net occurrence if it is not covered by another repeat, and the net frequency of a repeat is the number of its net occurrences. With this terminology, an NSMR is defined to be a repeat with a positive net frequency. Given the RLBWT of $T$, we show how to compute the set $\mathcal{S}^{\mathsf{nsmr}}$ of all NSMRs in $T$ together with their net frequency/occurrences in $O(n)$ time and $O(r)$ space. We also show that an $O(r)$-space data structure can be built from the RLBWT to support queries of computing the net frequency of any query pattern $P$ in optimal $O(|P|)$ time. The data structure is built in $O(r)$ space and in $O(n)$ time with high probability or deterministic $O(n+|\mathcal{S}^{\mathsf{nsmr}}|\log\log\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|))$ time, where $\sigma\leq r$ is the alphabet size of $T$. To achieve this, we prove that the total number of net occurrences is less than $2r$. With the duality between net occurrences and minimal unique substrings (MUSs), we get a new upper bound $2r$ of the number of MUSs in $T$, which may be of independent interest.
@@ 1 Introduction
Finding characteristic substring patterns in a given string $T$ is a fundamental task for string processing. One of the simplest patterns based on frequencies would be a repeat, which occurs at least twice in $T$. Since the number of distinct repeats in $T$ of length $n$ can be $\Theta(n^{2})$, maximality is often utilized to reduce the number down to $O(n)$ [13]. A substring of $T$ is called right-maximal (resp. left-maximal) if it cannot be extended to the right (resp. to the left) without reducing its number of occurrences in $T$. If a repeat is both right-maximal and left-maximal, it is called a maximal repeat (MR).
@@
Gusfield [13] also introduced stricter criteria to reduce the redundancy in maximal repeats. A repeat is called a supermaximal repeat (SMR) if it is not a substring of another repeat, and a near-supermaximal repeat (NSMR) if at least one of its occurrences is not covered by another repeat. It is easy to see that an SMR is an NSMR, and an NSMR is an MR. NSMRs were studied under the name of Chinese frequent strings in [17, 18, 28], and largest maximal repeats in [23, 7]. Gallé and Tealdi [8, 9] proposed the concept of context diversities to define a general class of context-sensitive repeats including MRs and SMRs.
@@
Enumerating these context-sensitive repeats has numerous applications in bioinformatics [13, 23, 1] and natural language processing [17, 18, 29, 20, 7] especially valuable for languages with no clear word segmentation such as Chinese and Japanese. Motivated by these applications, efficient enumeration algorithms have been extensively studied [29, 1, 27, 2, 26] using a family of suffix data structures such as suffix trees [31], directed acyclic word graphs (DAWGs) [5], compact DAWGs (CDAWGs) [4, 16], suffix arrays [19] and Burrows-Wheeler transform (BWT) [6].
@@
In this paper, we focus on r-enum [26], an algorithm to enumerate the MRs, minimal unique substrings (MUSs) [14] and minimal absent words (MAWs) in $O(r)$ words of compressed working space, where $r\leq n$ is the number of runs in the BWT of $T$ of length $n$. Given the run-length encoded BWT (RLBWT) of $T$, r-enum runs in $O(n\log\log_{w}(n/r))$ time in addition to the time linear to the number of output strings, where $w=\Theta(\log n)$ is the word size. We show that $O(n\log\log_{w}(n/r))$ time can be improved to $O(n)$ by resorting to the move data structure recently proposed by Nishimoto and Tabei [24] for constant-time LF-mapping on RLBWTs. We also extend r-enum to compute the NSMRs and SMRs, and the context diversity for every RMR in the same time and space complexities.
@@
Furthermore, we consider problems for net occurrences and net frequencies, originated from the studies of Chinese frequent strings [17, 18] and have recently attracted attention from combinatorial and algorithmic points of view [10, 28, 12, 15, 11, 21]. An occurrence of a repeat is called a net occurrence if the occurrence is not covered by another repeat, and the net frequency of a repeat is the number of its net occurrences. We would like to point out the fact, which may have been overlooked for over two decades, that the strings with positive net frequencies are equivalent to the NSMRs [13] and the largest maximal repeats [23, 7], and thus appear to have broader applications. Also, net occurrences are conceptually equivalent to maximum repeats in [14].
@@
Given the RLBWT of $T$, we show how to compute the set $\mathcal{S}^{\mathsf{nsmr}}$ of all NSMRs in $T$ together with their net frequency/occurrences in $O(n)$ time and $O(r)$ words of space, solving the task called All-NF in [10] on RLBWTs. We also show that an $O(r)$-space data structure can be built from the RLBWT to support queries of computing the net frequency of any query pattern $P$ in optimal $O(|P|)$ time. The data structure is built in $O(r)$ space and in $O(n)$ time with high probability or deterministic $O(n+|\mathcal{S}^{\mathsf{nsmr}}|\log\log\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|))$ time, where $\sigma\leq r$ is the alphabet size of $T$. To achieve this, we prove that the total number of net occurrences is less than $2r$. With the duality between net occurrences and MUSs shown in [14, 21], we immediately get a new upper bound $2r$ of the number of MUSs in $T$, which may be of independent interest.
@@ 2.2 Tools
Lemma 1 ([32, 30]).
@@
A predecessor query for an integer $i$ over the set $S$ of integers asks to compute $\max\{j\in S\mid j\leq i\}$, which can be efficiently supported by the data structure of [3, Appendix A]:
Lemma 2 ([3]).
@@
A range distinct query $\mathsf{RD}(x,p,q)$ for a string $x$ and integers $1\leq p\leq q\leq|x|$ asks to enumerate $(c,p_{c},q_{c})$ such that $p_{c}$ and respectively $q_{c}$ are the smallest and largest positions in $[b..e]$ with $x[p_{c}]=x[q_{c}]=c$ for every distinct character $c$ that occurs in $x[p..q]$. It is known that a range distinct query can be supported in output-optimal time after preprocessing $x$.
Lemma 3 ([22, 2]).
@@
A minor remark is that the output of range distinct queries of Lemma˜3 may not be sorted in any order of characters.
@@ 2.3 Context-sensitive repeats
To describe algorithms uniformly, we assume that the empty string is a repeat that occurs at every position $b$ in $[1..n]$, i.e., $\mathsf{Occ}(\varepsilon)=[1..n]$. Since $\mathsf{lc}(\varepsilon)=\mathsf{rc}(\varepsilon)=\Sigma_{\mathdollar}$, $\varepsilon$ is always in $\mathcal{S}^{\mathsf{lmr}}$, $\mathcal{S}^{\mathsf{rmr}}$ and $\mathcal{S}^{\mathsf{mr}}$. We set the following exception for the definition of net occurrences of $\varepsilon$: An occurrence $b\in\mathsf{Occ}(\varepsilon)$ is a net occurrence of $\varepsilon$ if and only if T[b1]T[b-1] and $T[b]$ are both unique. This exception provides a smooth transition to another characterization of net occurrences based on the uniqueness of extended net occurrences [10, 21].
@@ 2.4 Suffix trees and suffix arrays
The suffix trie of $T$ is the tree defined on the set $\mathcal{S}$ of nodes and edges from $x\in\mathcal{S}$ to $xc\in\mathcal{S}$ with $c\in\Sigma_{\mathdollar}$. The suffix tree of $T$ is the compacted suffix trie in which every non-branching internal node of the suffix trie is removed and considered to be an implicit node. Every internal node of the suffix tree is a right-maximal repeat $x$ and has $|\mathsf{rc}(x)|$ children. A link from a node $x$ to a (possibly implicit) node $ax\in\mathcal{S}$ for some $a\in\Sigma_{\mathdollar}$ is called a Weiner link. It is known that every node can be reached by following Weiner links from the root $\varepsilon$, and the total number of Weiner links is at most $3n$, which can be seen from the duality between suffix trees and DAWGs [5].

### v2

@@ Abstract
A repeat is a substring that occurs at least twice in a string, and is called a maximal repeat if it cannot be extended outwards without reducing its frequency. Nishimoto and Tabei [CPM, 2021] proposed r-enum, an algorithm to enumerate various characteristic substrings, including maximal repeats, in a string $T$ of length $n$ in $O(r)$ words of compressed working space, where $r\leq n$ is the number of runs in the Burrows-Wheeler transform (BWT) of $T$. Given the run-length encoded BWT (RLBWT) of $T$, r-enum runs in $O(n\log\log_{w}(n/r))$ time in addition to the time linear to the number of output strings, where $w=\Theta(\log n)$ is the word size. In this paper, we first improve the $O(n\log\log_{w}(n/r))$ term to $O(n)$. We next extend r-enum to compute other context-sensitive repeats such as near-supermaximal repeats (NSMRs) and supermaximal repeats, as well as the context diversity for every maximal repeat in the same complexities. Furthermore, we study net occurrences: An occurrence of a repeat is called a net occurrence if it is not covered by another repeat, and the net frequency of a repeat is the number of its net occurrences. With this terminology, an NSMR is a repeat with a positive net frequency. Given the RLBWT of $T$, we show how to compute the set $\mathcal{S}^{\mathsf{nsmr}}$ of all NSMRs in $T$ together with their net frequency/occurrences in $O(n)$ time and $O(r)$ space. We also show that an $O(r)$-space data structure can be built from the RLBWT to compute the net frequency/occurrences of any pattern in optimal time. The data structure is built in $O(r)$ space and in $O(n)$ time with high probability or deterministic $O(n+|\mathcal{S}^{\mathsf{nsmr}}|\log\log\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|))$ time, where $\sigma\leq r$ is the alphabet size of $T$. To achieve this, we prove that the total number of net occurrences is less than $2r$. With the duality between net occurrences and minimal unique substrings (MUSs), we get a new upper bound $2r$ of the number of MUSs in $T$, which may be of independent interest.
keywords:
Supermaximal repeats; Largest maximal repeats; Net frequencies; Run-length Burrows-Wheeler transform; Compressed data mining
@@ 1 Introduction
Finding characteristic substring patterns in a given string $T$ is a fundamental task for string processing. One of the simplest patterns based on frequencies would be a repeat, which occurs at least twice in $T$. Since the number of distinct repeats in $T$ of length $n$ can be $\Theta(n^{2})$, maximality is often utilized to reduce the number down to $O(n)$ [1997Gusfield_AlgorOnStrinTreesAnd]. A substring of $T$ is called a right-maximal repeat (RMR) (resp. left-maximal repeat (LMR)) if it cannot be extended to the right (resp. to the left) without reducing its number of occurrences in $T$. If a repeat is both right-maximal and left-maximal, it is called a maximal repeat (MR).
@@
Gusfield [1997Gusfield_AlgorOnStrinTreesAnd] also introduced stricter criteria to reduce the redundancy in maximal repeats. A repeat is called a supermaximal repeat (SMR) if it is not a substring of another repeat, and a near-supermaximal repeat (NSMR) if at least one of its occurrences is not covered by another repeat. It is easy to see that an SMR is an NSMR, and an NSMR is an MR. NSMRs were studied under the name of Chinese frequent strings in [2001LinY_ExtracChinesFrequenStrinWithout, 2004LinY_ProperAndFurthApplicOf, 2024OhlebuschBO_FasterComputOfChinesFrequen_SPIRE], and largest maximal repeats in [2008NicolasRSPCDTVM_ModelLocalRepeatOnGenom, 2013Galle_BagOfRepeatRepresOf_SIGIR]. Gallé and Tealdi [2014GalleT_ContexDiverRepeatAndTheir_LATA, 2018GalleT_EmphxRepeatNewTaxonOf] proposed the concept of context diversities to define a general class of context-sensitive repeats including MRs and SMRs.
@@
Enumerating these context-sensitive repeats has numerous applications in bioinformatics [1997Gusfield_AlgorOnStrinTreesAnd, 2008NicolasRSPCDTVM_ModelLocalRepeatOnGenom, 2009BecherDH_EfficComputOfAllPerfec] and natural language processing [2001LinY_ExtracChinesFrequenStrinWithout, 2004LinY_ProperAndFurthApplicOf, 2009OkanoharaT_TextCategWithAllSubst_SDM, 2011MasadaTSO_ClustDocumWithMaximSubst, 2013Galle_BagOfRepeatRepresOf_SIGIR] especially valuable for languages with no clear word segmentation such as Chinese and Japanese. Motivated by these applications, efficient enumeration algorithms have been extensively studied [2009OkanoharaT_TextCategWithAllSubst_SDM, 2009BecherDH_EfficComputOfAllPerfec, 2015OhlebuschB_AlphabIndepAlgorForFindin, 2020BelazzouguiCKM_LinearTimeStrinIndexAnd, 2021NishimotoT_REnumEnumerOfCharac_CPM] using a family of suffix data structures such as suffix trees [1973Weiner_LinearPatterMatchAlgor], directed acyclic word graphs (DAWGs) [1985BlumerBHECS_SmallAutomRecogSubworOf], compact DAWGs (CDAWGs) [1987BlumerBHME_ComplInverFilesForEffic, 2005InenagaHSTAMP_LineConstOfCompacDirec], suffix arrays [1993ManberM_SuffixArrayNewMethodFor, JDA2004AbouelhodaKO_ReplacSuffixTreesWithEnhan] and Burrows-Wheeler transform (BWT) [1994BurrowsW_BlockSortinLosslDataCompr].
@@
In this paper, we focus on r-enum [2021NishimotoT_REnumEnumerOfCharac_CPM], an algorithm to enumerate the MRs, minimal unique substrings (MUSs) [2011IlieS_MinimUniqueSubstAndMaxim] and minimal absent words (MAWs) in $O(r)$ words of compressed working space, where $r\leq n$ is the number of runs in the BWT of $T$ of length $n$. Given the run-length encoded BWT (RLBWT) of $T$, r-enum runs in $O(n\log\log_{w}(n/r))$ time in addition to the time linear to the number of output strings, where $w=\Theta(\log n)$ is the word size. We show that $O(n\log\log_{w}(n/r))$ time can be improved to $O(n)$ by resorting to the move data structure recently proposed by Nishimoto and Tabei [2021NishimotoT_OptimTimeQueriesOnBwt_ICALP] for constant-time LF-mapping on RLBWTs. We also extend r-enum to compute the NSMRs and SMRs, and the context diversity for every RMR in the same time and space complexities.
@@
Furthermore, we consider problems for net occurrences and net frequencies, originated from the studies of Chinese frequent strings [2001LinY_ExtracChinesFrequenStrinWithout, 2004LinY_ProperAndFurthApplicOf] and have recently attracted attention from combinatorial and algorithmic points of view [2024GuoEWZ_ExploitNewProperOfStrin_CPM, 2024OhlebuschBO_FasterComputOfChinesFrequen_SPIRE, 2024GuoUWZ_OnlinComputOfStrinNet_SPIRE, 2024Inenaga_FasterAndSimplOnlinComput, 2025GuoK_NetOccurInFibonAnd_CPM, 2025MienoI_SpaceEfficOnlinComputOf_CPM]. An occurrence of a repeat is called a net occurrence if the occurrence is not covered by another repeat, and the net frequency of a repeat is the number of its net occurrences. We would like to point out the fact, which may have been overlooked for over two decades, that the strings with positive net frequencies are equivalent to the NSMRs [1997Gusfield_AlgorOnStrinTreesAnd] and the largest maximal repeats [2008NicolasRSPCDTVM_ModelLocalRepeatOnGenom, 2013Galle_BagOfRepeatRepresOf_SIGIR], and thus appear to have broader applications. Also, net occurrences are conceptually equivalent to maximum repeats in [2011IlieS_MinimUniqueSubstAndMaxim].
@@
Given the RLBWT of $T$, we show how to compute the set $\mathcal{S}^{\mathsf{nsmr}}$ of all NSMRs in $T$ together with their net frequency/occurrences in $O(n)$ time and $O(r)$ words of space, solving the task called All-NF in [2024GuoEWZ_ExploitNewProperOfStrin_CPM] on RLBWTs. We also address the task called Single-NF [2024GuoEWZ_ExploitNewProperOfStrin_CPM], which requires building a data structure to support NF-queries of computing the net frequency and/or net occurrences of any query pattern. We show that an $O(r)$-space data structure can be built from the RLBWT to support NF-queries in optimal time. The data structure is built in $O(r)$ space and in $O(n)$ time with high probability or deterministic $O(n+|\mathcal{S}^{\mathsf{nsmr}}|\log\log\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|))$ time, where $\sigma\leq r$ is the alphabet size of $T$. To achieve this, we prove that the total number of net occurrences is less than $2r$. With the duality between net occurrences and MUSs shown in [2011IlieS_MinimUniqueSubstAndMaxim, 2025MienoI_SpaceEfficOnlinComputOf_CPM], we immediately get a new upper bound $2r$ of the number of MUSs in $T$, which may be of independent interest.
1.1 Related work
Context-sensitive repeats are closely related to and efficiently computed from suffix data structures. A folklore algorithm [1997Gusfield_AlgorOnStrinTreesAnd] uses suffix trees. To reduce the (rather big) working space of suffix trees by a constant factor, algorithms based on enhanced suffix arrays [JDA2004AbouelhodaKO_ReplacSuffixTreesWithEnhan] have been proposed. A compact- or compressed-space solution can be achieved by using LF-mapping and range distinct queries (see Section˜2) on BWTs or RLBWTs to simulate a traversal on suffix link trees. Table˜1 summarizes representative algorithms.
reference:
used techniques
space (bits) running time patterns
[1997Gusfield_AlgorOnStrinTreesAnd]: suffix tree $O(n\log n)$ $O(n\log\sigma)$ MR, NSMR, SMR
[2009OkanoharaT_TextCategWithAllSubst_SDM]: enhanced suffix array $O(n\log n)$ $O(n)$ MR
[2015OhlebuschB_AlphabIndepAlgorForFindin]: enhanced suffix array $O(n\log n)$ $O(n)$
MR, NSMR, SMR,
context diversity
[2020BelazzouguiCKM_LinearTimeStrinIndexAnd]: BWT and RD $O(n\log\sigma)$ $O(n)$ MR, MUS, MAW
[2012BellerBO_SpaceEfficComputOfMaxim_SPIRE]: BWT and RD $|\mathit{RD}|+O(n)$ $O(nd)$ MR, SMR
[2015BelazzouguiC_SpaceEfficDetecOfUnusual_SPIRE]: BWT and RD $|\mathit{RD}|+O(\sigma^{2}\log^{2}n)$ $O(nd)$ MR, MUS, MAW
[2021NishimotoT_REnumEnumerOfCharac_CPM]: RLBWT and RD $O(r\log n)$ $O(n\log\log_{w}(n/r))$ MR, MUS, MAW
this work:
RLBWT and RD
$O(r\log n)$ $O(n)$
MR, MUS, MAW,
NSMR, SMR,
context diversity
Table 1: Comparison between algorithms that find characteristic substring patterns in a string $T$ of length $n$ over an integer alphabet of size $\sigma=n^{O(1)}$, where $w=\Theta(\log n)$ is the word size. Each method takes $T$, its BWT or $O(r)$-size RLBWT as an input, and enumerate substring patterns represented in $O(1)$ space each. For MAWs, we additionally need the time linear to the number of output, which is bounded by $O(n\sigma)$. RD represents a data structure for range distinct queries, which can be implemented in $|\mathit{RD}|$ bits of space while supporting a query in $O(d)$ time per output with $(|\mathit{RD}|,d)=(n\log_{2}\sigma+o(n\log\sigma),O(\log\sigma))$ using Lemma˜2.3 or $(|\mathit{RD}|,d)=(n\log_{2}\sigma+o(n\log\sigma),O(\log\sigma))$ [2015ClaudeNP_WavelMatrixEfficWavelTree].
The study for net frequencies/occurrences was initiated recently by Guo et al. [2024GuoEWZ_ExploitNewProperOfStrin_CPM]. Given a string $T$ of length $n$ over an integer alphabet of size $\sigma=n^{O(1)}$, we can perform All-NF in $O(n)$ time and words of space using an algorithm [2015OhlebuschB_AlphabIndepAlgorForFindin, 2024OhlebuschBO_FasterComputOfChinesFrequen_SPIRE] based on enhanced suffix arrays. The problem has also been studied in an online setting, where we maintain a data structure to perform All-NF and Single-NF while allowing monotonic updates to the left and/or right ends of $T$. The current state-of-the-art online algorithm [2025MienoI_SpaceEfficOnlinComputOf_CPM] is based on online construction of implicit suffix trees and implicit CDAWGs. In this paper, we achieve the first offline algorithm for All-NF and Single-NF that works in $O(r)$ words of compressed space.
1.2 Organization of paper
This paper is organized as follows. In Section˜2, we introduce notation and tools we use in the following sections. In Section˜3, we prove that the total number of net occurrences is less than $2r$, which is crucial to bound the time and space complexities of our algorithms for net frequencies/occurrences. In Section˜4, we speed up r-enum and extend it to enumerate other context-sensitive repeats. In Section˜5, we present an $O(r)$-space data structure for NF-queries based on RLBWTs. In Section˜6, we conclude the paper with some remarks.
@@ 2.2 Tools
Lemma 2.1 ([2000Willard_ExaminComputGeometVanEmde, 2008Ruzic_ConstEfficDictionInClose_ICALP]).
@@
A predecessor query for an integer $i$ over the set $S$ of integers asks to compute $\max\{j\in S\mid j\leq i\}$, which can be efficiently supported by the data structure of [2015BelazzouguiN_OptimLowerAndUpperBound, Appendix A]:
Lemma 2.2 ([2015BelazzouguiN_OptimLowerAndUpperBound]).
@@
A range distinct query $\mathsf{RD}(x,p,q)$ for a string $x$ and integers $1\leq p\leq q\leq|x|$ asks to enumerate $(c,p_{c},q_{c})$ such that $p_{c}$ and respectively $q_{c}$ are the smallest and largest positions in $[p..q]$ with $x[p_{c}]=x[q_{c}]=c$ for every distinct character $c$ that occurs in $x[p..q]$. It is known that a range distinct query can be supported in output-optimal time after preprocessing $x$.
Lemma 2.3 ([2002Muthukrishnan_EfficAlgorForDocumRetriev_SODA, 2020BelazzouguiCKM_LinearTimeStrinIndexAnd]).
@@
A minor remark is that the output of range distinct queries of Lemma˜2.3 may not be sorted in any order of characters.
@@ 2.3 Context-sensitive repeats
To describe algorithms uniformly, we assume that the empty string is a repeat that occurs at every position $b$ in $[1..n]$, i.e., $\mathsf{Occ}(\varepsilon)=[1..n]$. Since $\mathsf{lc}(\varepsilon)=\mathsf{rc}(\varepsilon)=\Sigma_{\mathdollar}$, $\varepsilon$ is always in $\mathcal{S}^{\mathsf{lmr}}$, $\mathcal{S}^{\mathsf{rmr}}$ and $\mathcal{S}^{\mathsf{mr}}$. We set the following exception for the definition of net occurrences of $\varepsilon$: An occurrence $b\in\mathsf{Occ}(\varepsilon)$ is a net occurrence of $\varepsilon$ if and only if T[b1]T[b-1] and $T[b]$ are both unique. This exception provides a smooth transition to another characterization of net occurrences based on the uniqueness of extended net occurrences [2024GuoEWZ_ExploitNewProperOfStrin_CPM, 2025MienoI_SpaceEfficOnlinComputOf_CPM].
@@ 2.4 Suffix trees and suffix arrays
The suffix trie of $T$ is the tree defined on the set $\mathcal{S}$ of nodes and edges from $x\in\mathcal{S}$ to $xc\in\mathcal{S}$ with $c\in\Sigma_{\mathdollar}$. The suffix tree of $T$ is the compacted suffix trie in which every non-branching internal node of the suffix trie is removed and considered to be an implicit node. Every internal node of the suffix tree is a right-maximal repeat $x$ and has $|\mathsf{rc}(x)|$ children. A link from a node $x$ to a (possibly implicit) node $ax\in\mathcal{S}$ for some $a\in\Sigma_{\mathdollar}$ is called a Weiner link. It is known that every node can be reached by following Weiner links from the root $\varepsilon$, and the total number of Weiner links is at most $3n$, which can be seen from the duality between suffix trees and DAWGs [1985BlumerBHECS_SmallAutomRecogSubworOf].